LuciaA24:

The cost of fabric varies directly with the amount of fabric purchased. If 123 yards of fabric cost $10.00, how much would 412 yards cost?

sllo:

123 plus 10 bucks 10123 is your answer

GoatyGoat:

$33.51 is the answer

GoatyGoat:

totally didn't use google right there

Lilly2010:

lol

toga:

We can use the concept of direct variation to solve this problem. Direct variation is a relationship between two variables that can be expressed as y = kx, where y and x are the variables, and k is a constant of proportionality. In this problem, the amount of fabric purchased (x) is directly proportional to the cost of fabric (y). So, we can write: y = kx To find the value of k, we can use the information given in the problem. When x = 123, y = $10.00. So: 10 = k * 123 Solving for k, we get: k = 10/123 Now we can use this value of k to find the cost of 412 yards of fabric: y = (10/123) * 412 y = $33.50 Therefore, 412 yards of fabric would cost $33.50.

GoatyGoat:

OH I was 1 cent off

toga:

@goatygoat wrote:
OH I was 1 cent off
its just one cent no big deal

GoatyGoat:

@toga wrote:
@goatygoat wrote:
OH I was 1 cent off
its just one cent no big deal
Ok